two years ago,Perla was 7 times as old as Merly.two years from now,Perla will be 6 years more than 4 times as old as merly will be.Find the present age of each.
Answer by TimothyLamb(4379)

---
x = Perla
y = Merly
---
x - 2 = 7(y - 2)
x + 2 = 4(y + 2) + 6
---
put the system of linear equations into standard form:
---
x - 2 = 7y - 14
x + 2 = 4y + 8 + 6
---
x - 7y = -12
x - 4y = 12
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = Perla = 44
y = Merly = 8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php